Differential Analysis for a Discontinued Product

A condensed income statement by product line for Warrick Beverage Inc. indicated the following for Mango Cola for the past year:

Sales $15,000,000
Cost of goods sold (10,800,000)
Gross profit $4,200,000
Operating expenses (8,000,000)
Operating loss $(3,800,000)

It is estimated that 30% of the cost of goods sold represents fixed factory overhead costs and that 25% of the operating expenses are fixed. Because Mango Cola is only one of many products, the fixed costs will not be materially affected if the product is discontinued.

a. Prepare a differential analysis dated February 29 to determine whether Mango Cola should be continued (Alternative 1) or discontinued (Alternative 2). If an amount is zero, enter "0". If required, use a minus sign to indicate a loss.

Differential Analysis
Continue (Alt. 1) or Discontinue (Alt. 2) Mango Cola
February 29
Continue
Mango Cola
(Alternative 1)
Discontinue
Mango Cola
(Alternative 2)
Differential
Effects
(Alternative 2)
Revenues $ $ $
Costs:
Variable cost of goods sold
Variable operating expenses
Fixed costs
Profit (Loss) $ $ $
Answer #1

Answer-

WARRICK BEVERAGE INC.
DIFFERENTIAL ANALYSIS
ALTERNATIVE 1 ALTERNATIVE 2 DIFFERENTIAL EFFECT
CONTINUE MANGO COLA DISCONTINUE MANGO COLA
PARTICULARS $ $ $
Sales 15000000 0 -15000000
Less- Variable costs
Cost of goods sold $10800000*70%= $7560000 0 7560000
Operating expenses $8000000*75%= $6000000 0 6000000
Contribution margin 1440000 -1440000
Less- Fixed costs
Factory overhead $10800000*30%= $3240000 3240000 0
Operating expenses $8000000*25%= $2000000 2000000 0
Net operating income (loss) -3800000 -5240000 -1440000

Hence the company should not discontinue its mango cola product line otherwise the company will loss it profit by $1440000.